What can be the result of adding this 2 equations?


4x+y=7
X-y=-2

Answer by ewatrrr(24785)


Hi
by adding...we find the point the two lines may have in common.
4x+y= 7
x-y=-2
5x = 5
x = 1
4x+y= 7
4*1+y= 7
y = 3
Lines intersect at Pt(1,3)which is the solution for the system